# SpikeNAS: A Fast Memory-Aware Neural Architecture Search Framework for Spiking Neural Network-based Autonomous Agents

### v3

@@ Abstract
Autonomous mobile agents (e.g., UAVs and UGVs) are typically expected to incur low power/energy consumption for solving machine learning tasks (such as object recognition), as these mobile agents are usually powered by portable batteries. These requirements can be fulfilled by Spiking Neural Networks (SNNs), since their bio-inspired spike-based operations offer high accuracy and ultra low-power/energy computation. Currently, most of the SNN architectures are derived from Artificial Neural Networks whose neurons’ architectures and operations are different from SNNs, or developed without considering memory budgets from the underlying processing hardware of autonomous mobile agents. These limitations hinder SNNs from reaching their full potential in accuracy and efficiency. Toward this, we propose SpikeNAS, a novel fast memory-aware neural architecture search (NAS) framework for SNNs that quickly finds an appropriate SNN architecture with high accuracy under the given memory budgets from autonomous mobile agents. To do this, our SpikeNAS employs several key steps: analyzing the impacts of network operations on the accuracy, enhancing the network architecture to improve the learning quality, and developing a fast memory-aware search algorithm. The experimental results show that our SpikeNAS improves the searching time and maintains high accuracy as compared to state-of-the-art while meeting the given memory budgets (e.g., 4.4x faster search with 1.3% accuracy improvement for CIFAR100, using an Nvidia RTX 6000 Ada GPU machine), thereby quickly providing the appropriate SNN architecture for the memory-constrained autonomous mobile agents.
I Introduction
In recent years, autonomous mobile agents (e.g., UAVs, UGVs, and mobile robots) are expected to incorporate neural network (NN) algorithms [1, 2, 3], as these algorithms achieve state-of-art accuracy when solving diverse machine learning (ML) tasks, such as image classification [4], object recognition [5], and continual learning [6]. These mobile agents are also expected to incur low power/energy consumption when executing NN algorithms, because they are usually powered by portable batteries [2][3]. These requirements can be fulfilled by Spiking Neural Networks (SNNs), as their bio-inspired spike-based operations offer high accuracy while incurring ultra low-power/energy computation [7]. Therefore, many research works are carried out to maximize the SNN potentials, ranging from SNN architecture developments to its deployments for diverse ML-based applications [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. In fact, different ML-based applications usually have different accuracy requirements, which often necessitate different memory costs to facilitate the corresponding SNN architectures. For instance, a bigger SNN is typically developed to achieve a higher accuracy, but at the cost of a larger memory footprint, as shown in Fig. 1(a). Furthermore, a larger memory footprint leads to a higher energy consumption, as the memory accesses still dominate the overall processing energy of SNNs; see Fig. 1(b). Hence, the development of SNN architectures that can achieve the expected accuracy with acceptable memory footprint is needed to support the widespread use of SNNs for autonomous agents.
@@ I Introduction
Figure 1: (a) Autonomous mobile agents are typically expected to achieve high accuracy with low power and energy-efficient processing to complete ML-based tasks. (b) Accuracy and memory of different SNN models for CIFAR10: CIFARNet1 [18], CIFARNet2 [19], ResNet11 [20], ResNet19 [21], AutoSNN [13]; adapted from studies in [13]. (c) Breakdown of energy consumption for SNN processing by different hardware platforms: PEASE [22], TrueNorth [23], and SNNAP [24]; adapted from studies in [25].
@@
Currently, most of the SNN architectures are derived from Artificial Neural Networks (ANNs), whose neurons’ architectures and operations are different from SNNs [13]. SNNs accumulate and generate temporal information in form of spikes using Leaky Integrate-and-Fire (LIF) neurons, while ANNs perform multiplication-and-accumulation (MAC) and activation using Rectified Linear Unit (ReLU). Such ANN-based SNN architectures may lose unique features from SNN computation models (e.g., temporal information between spikes), thereby leading to sub-optimal accuracy [14]. Moreover, the SNN architectures are still developed without considering memory budgets from the underlying processing hardware, which is important to ensure the applicability of the SNN architecture for the given hardware platform. For instance, most of the SNN models require more than 20M of parameters to achieve 93% accuracy for CIFAR10, as shown in Fig. 1(a), which is memory consuming and inefficient for resource-constrained autonomous mobile agents. These limitations hinder the SNNs from reaching their full potential in accuracy and efficiency for autonomous mobile agents. However, manually designing the desired SNN architecture is a laborious task and time consuming, thus requiring an alternative solution that quickly provides an SNN architecture that is suitable for the given accuracy and memory constraints.
@@
Research Problem: How to automatically and quickly find an SNN architecture that achieves acceptable accuracy under the given memory constraints for autonomous mobile agents. An efficient solution to this problem will ease SNN developments for autonomous mobile agents, thereby enabling the widespread use of SNNs for diverse ML-based tasks [26].
I-A State-of-the-Art and Their Limitations
To address the research problem, previous works employ neural architecture search (NAS) to automatically find SNN architectures that can achieve high accuracy. Currently, NAS for SNNs is still at the early stage, hence only a few studies found in the literature [13][14][27, 28, 29], and the prominent ones are discussed in the following.
  • AutoSNN [13]: It aims at finding an SNN architecture with high accuracy and low number of spikes by employing a NAS with training approach. First, it trains a super-network that includes all of the possible candidates of SNN architectures, then performs a search to find a subset of network architecture that achieves the target. However, this NAS approach requires a huge searching time as well as a huge memory and energy consumption, as it needs to perform training for a large network before searching and determining the desired SNN architecture.
  • SNASNet [14]: It aims at finding an SNN architecture that achieves high accuracy by employing a NAS without training approach to reduce the searching and training time. It is performed by exploring the architecture candidates of the cell (i.e., a directed acyclic graph that represents network topology and operations; see Fig. 2a), evaluating the score of the investigated architecture, then training the architecture with the highest score [14]. However, it has several limitations as follows.
    • It has a fixed number of cells (i.e., 2 cells), and it applies the same architecture for all cells, thereby reducing the possibility to find a better SNN architecture (e.g., higher accuracy).
Moreover, these state-of-the-art works also do not consider memory constraints from the target hardware in their search algorithms, thereby limiting the efficiency of the generated SNN architectures when deployed on the target hardware. To identify and better understand the limitations of state-of-the-art (i.e., a NAS without training approach), we perform an experimental case study and discuss it in Section I-B.
I-B Case Study and Research Challenges
@@
We study the impacts of search iterations and operation types on the accuracy considering a NAS without training approach using the search space and mechanism of SNASNet, i.e, exploring the topology and operations of a cell architecture in Fig. 2(a) using a random search. To evaluate the impact of search iterations, the state-of-the-art work employs 5000x iterations and the investigated one employs 4096x. To evaluate the impact of operation types, the state-of-the-art work employs 5 operation types in Fig. 2(a), while the investigated one employs 4 operation types (i.e., without Zeroize). Details of the experimental setup are provided in Section IV. The experimental results are presented in Fig. 2(b), from which we draw the following key observations.
@@
  • The solution should incorporate the memory constraints into its searching algorithm to ensure the applicability and efficiency of the SNN deployments on the targeted hardware platforms.
I-C Our Novel Contributions
Towards this, we propose SpikeNAS, a novel NAS framework that can quickly find the appropriate SNN architecture with high accuracy under the given memory budgets from autonomous mobile agents. Our SpikeNAS employs several key steps as the following (see an overview in Fig. 3, and the details in Fig. 4).
  • Analysis of the Impacts of Network Operations: It aims to understand the significance of each pre-defined operation type. It is done by removing the investigated operation type from the search space and then observing the output accuracy.
  • Network Architecture Enhancements: It aims to maintain (or improve) accuracy of the network architecture, while reducing the searching time. It is done by optimizing the cell operation types and the number of cells in the network architectures.
  • A Fast Memory-aware Search Algorithm: It aims to find an SNN architecture that achieves high accuracy and meets the memory budget, by performing search for each individual cell, while monitoring the memory cost of the investigated architecture.
Key Results: We evaluate our SpikeNAS framework using a Python-based implementation on an Nvidia RTX 6000 Ada GPU machine. The experimental results show that our SpikeNAS improves the searching time and maintains high accuracy as compared to state-of-the-art while meeting the memory budgets, e.g., SpikeNAS improves the searching time by 4.4x and accuracy by 1.3% under 2M parameters for CIFAR100.
@@ I-C Our Novel Contributions
Figure 4: Overview of our SpikeNAS framework, and its key steps for generating an SNN architecture that offers high accuracy while meeting the given memory constraint from an autonomous mobile agent. The novel contributions are highlighted in blue boxes.
II Preliminaries
II-A Spiking Neural Networks (SNNs)
SNNs are considered the third generation of NN computation model [31]. SNNs typically employ a LIF neuron model [32, 33, 34, 35, 36], whose neuronal dynamics can be formulated as follows.
@@ II-A Spiking Neural Networks (SNNs)
II-B NAS without Training
NAS without training approach was proposed in ANN domain [37], and then leveraged for SNN domain [14]. Its key idea is to quickly measure the capability of a given network to reach high accuracy at early stage, thereby avoiding the training phase to evaluate the network. It is motivated by the observation that an architecture with distinctive representations across different samples is likely to achieve high accuracy after training [37]. To measure the representation capability of a network, the state-of-the-art [14] employs the following key steps (our SpikeNAS framework also follows the same steps for evaluating the network).
@@ II-B NAS without Training
II-C Neural Cell-based Strategy for NAS
Neural cell (or simply cell)-based strategy has been widely used for NAS in ANNs. The motivation of cell-based NAS is to provide a unified benchmark for NAS algorithms [30]. Its approach is to search for the connection topology and the corresponding operations inside each cell; see Fig. 2(a). A cell typically consists of 4 nodes and 5 pre-defined operation types: No-Connection (Zeroize), Skip-Connection (SkipCon), 1x1 Convolution (1x1Conv), 3x3 Convolution (3x3Conv), and 3x3 Average Pooling (3x3AvgPool). The connection between 2 nodes employs a specific pre-defined operation type. Therefore, there are 15625 architecture candidates in each cell. The state-of-the-art (i.e., SNASNet) employs a 2-cell SNN architecture for its NAS technique; see its macro-architecture in Fig. 5. In this work, we also employ the same macro-architecture for our SpikeNAS framework, and focus on the feedforward topology as it is widely used in the SNN community.
III SpikeNAS Framework
To address the above-discussed research problem and challenges, we propose a novel SpikeNAS framework; see the detailed flow in Fig. 4. Its key steps are discussed in the Section III-A - Section III-C, consecutively.
III-A Analysis of the Impact of Operation Types
Discussion in Section I-B suggests that employing all pre-defined operations might not lead to optimal solution. Therefore, to understand the significance of each pre-defined operation, we perform an experimental case study. Here, we remove a specific operation type, then perform NAS considering a macro-architecture in Fig. 5 with 5000x search iteration. Since we employ the NAS without training approach, then this step is performed before the search process and the training phase. Experimental results are shown in Fig. 6, from which we draw the following key observations.
@@ III-A Analysis of the Impact of Operation Types
    Removing the 3x3Conv operation from the cells significantly reduces the accuracy, as this operation employs 3x3 kernel size for extracting important features from input feature maps. Therefore, removing the 3x3Conv make the network difficult to perform learning properly. Therefore, the 3x3Conv operation should be kept in the search space.
@@
III-B Network Architecture Enhancements
III-B1 Optimization of the Operation Types in the Neural Cells
In Section III-A, we observe that some operations can be removed from the search space to improve the searching time without sacrificing accuracy. Hence, we optimize the number of operation types based on their significance on the accuracy. We propose two variants of operation sets that can be considered in the search space to provide multiple optimization options (see an overview in Fig. 7), as follows.
@@ III-B1 Optimization of the Operation Types in the Neural Cells
III-B2 SNN Macro-Architecture Adjustments
@@
  • A 1-Cell Architecture: It aims at reducing the memory footprint, but might reduce the accuracy as compared to the 2-cell architecture; see an illustration in Fig. 8. Its key idea is removing one cell from the 2-cell architecture; see an overview in Fig. 9(a).
  • A 3-Cell Architecture: It aims at improving the accuracy, but might incur a larger memory footprint as compared to the 2-cell architecture; see an illustration in Fig. 8. Its key idea is adding one cell to the 2-cell architecture; see an overview in Fig. 9(b).
@@
III-C A Fast Memory-aware Search Algorithm

### v4

# SpikeNAS: A Fast Memory-Aware Neural Architecture Search Framework for Spiking Neural Network-based Embedded AI Systems

@@ Abstract
Embedded AI systems are expected to incur low power/energy consumption for solving machine learning tasks, as these systems are usually power constrained (e.g., object recognition task in autonomous mobile agents with portable batteries). These requirements can be fulfilled by Spiking Neural Networks (SNNs), since their bio-inspired spike-based operations offer high accuracy and ultra low-power/energy computation. Currently, most of SNN architectures are derived from Artificial Neural Networks whose neurons’ architectures and operations are different from SNNs, and/or developed without considering memory budgets from the underlying processing hardware of embedded platforms. These limitations hinder SNNs from reaching their full potential in accuracy and efficiency. Toward this, we propose SpikeNAS, a novel fast memory-aware neural architecture search (NAS) framework for SNNs that quickly finds an appropriate SNN architecture with high accuracy under the given memory budgets from targeted embedded systems. To do this, our SpikeNAS employs several key steps: analyzing the impacts of network operations on the accuracy, enhancing the network architecture to improve the learning quality, developing a fast memory-aware search algorithm, and performing quantization. The experimental results show that our SpikeNAS improves the searching time and maintains high accuracy compared to state-of-the-art while meeting the given memory budgets (e.g., 29x, 117x, and 3.7x faster search for CIFAR10, CIFAR100, and TinyImageNet200 respectively, using an Nvidia RTX A6000 GPU machine), thereby quickly providing the appropriate SNN architecture for the memory-constrained embedded AI systems.
{IEEEImpStatement}
Developing SNN models for the memory- and power-constrained embedded AI systems is highly desirable due to its immense benefits, but very challenging to do. However, manual SNN development is laborious and time-consuming, while state-of-the-art NAS techniques for SNNs do not consider the memory budget from the targeted hardware platform. To address this, our NAS technique quickly generates an SNN that meets memory budget and offers high accuracy by identifying the impact of network operations, enhancing the network architecture, and developing a fast search algorithm. In this manner, our research contributes to the development of application-specific SNNs under any given hardware platforms, thereby enabling their applicability for diverse application use-cases.
{IEEEkeywords}
Embedded AI systems, fast and memory-aware, neural architecture search (NAS), network enhancements, optimizations, spiking neural networks (SNNs).
1 Introduction
@@
Figure 1: (a) Embedded AI systems are expected to achieve high accuracy with low power and energy-efficient processing to complete ML-based tasks. (b) Accuracy and memory of different SNN models for CIFAR10: CIFARNet1 [1], CIFARNet2 [2], ResNet11 [3], ResNet19 [4], AutoSNN [5]; adapted from studies in [5]. (c) Breakdown of energy consumption for SNN processing by different hardware platforms: PEASE [6], TrueNorth [7], and SNNAP [8]; adapted from studies in [9].
[…]
Emdedded systems are expected to incorporate neural network (NN) algorithms [10, 11, 12, 13], since these algorithms can achieve state-of-art accuracy in solving diverse machine learning (ML) tasks, such as image classification [14], object recognition [15], and continual learning [16]; referred to as the embedded artificial intelligence (AI) systems. Therefore, these systems need to incur low power and energy-efficient consumption when executing NN algorithms, because they are usually powered by portable batteries [11][12]. For instance, battery-powered autonomous mobile agents (e.g., UGVs) need to achieve high accuracy with efficient NN processing to preserve their battery lifetime; see Fig. 1(a). These requirements can be fulfilled by Spiking Neural Networks (SNNs), as their bio-inspired spike-based operations offer high accuracy while incurring ultra low-power/energy computation [17]. Therefore, many research works are carried out to maximize the SNN potentials, ranging from SNN architecture developments to its deployments for diverse ML-based applications [5][18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. In fact, different ML-based applications usually have different accuracy requirements, which often necessitate different memory costs to facilitate the corresponding SNN architectures. For instance, a bigger SNN is typically developed to achieve a higher accuracy, while incurring a larger memory footprint, as shown in Fig. 1(b). Furthermore, a larger memory footprint leads to a higher energy consumption, as the memory accesses still dominate the overall processing energy of SNNs; see Fig. 1(c). Therefore, the development of SNN architectures that can achieve the expected accuracy with acceptable memory footprint is needed to support the widespread use of SNNs for embedded AI systems.
@@
Currently, most of the SNN architectures are derived from Artificial Neural Networks (ANNs), whose neurons’ architectures and operations are different from SNNs [5]. SNNs accumulate and generate temporal information in form of spikes using Leaky Integrate-and-Fire (LIF) neurons, while ANNs perform multiplication-and-accumulation (MAC) and activation using Rectified Linear Unit (ReLU). Such ANN-based SNN architectures may lose unique features from SNN computation models (e.g., temporal information between spikes), thereby leading to sub-optimal accuracy [28]. Moreover, the SNN architectures are still developed without considering memory budgets from the underlying processing hardware, which is important to ensure the applicability of the SNN architecture for the given hardware platform. For instance, most of the SNN models require more than 2x107 (20M) of parameters to achieve 93% accuracy for the CIFAR10 dataset, as shown in Fig. 1(b), which is memory consuming and inefficient for resource-constrained embedded systems. These limitations hinder the SNNs from reaching their full potential in accuracy and efficiency for embedded AI systems. However, manually designing the desired SNN architecture is a laborious task and time consuming, thus requiring an alternative solution that quickly provides an SNN architecture that is suitable for the given accuracy and memory constraints.
@@
Research Problem: How to automatically and quickly find an SNN architecture that achieves acceptable accuracy under the given memory constraints for embedded AI systems. An efficient solution to this problem will ease SNN developments for embedded AI systems, thus enabling the widespread use of SNNs for diverse ML-based application use-cases [13].
@@
1.1 State-of-the-Art and Their Limitations
To address the research problem, previous works employ neural architecture search (NAS) to automatically find SNN architectures that can achieve high accuracy. Currently, NAS for SNNs is still at the early stage, hence only a few studies found in the literature, including [5][28][39, 40, 41]. The prominent one is SNASNet [28], which finds an SNN architecture that achieves high accuracy by employing a NAS without training approach to reduce the searching time. However, these state-of-the-art works do not consider memory constraints from the target hardware in their search algorithms, thereby limiting the efficiency of the generated SNN architectures when deployed on the target hardware. To identify and better understand the limitations of state-of-the-art (i.e., the NAS without training approach in [28]), we perform an experimental case study and discuss it in Section 1.2.
1.2 Case Study and Research Challenges
We study the impacts of search iterations and operation types on the accuracy considering a NAS without training approach using the search space and mechanism of SNASNet, i.e, exploring the topology and operations of a cell architecture in Fig. 2(a) using a random search. To evaluate the impact of search iterations, the state-of-the-art work employs 5000x iterations and the investigated one employs 4096x. To evaluate the impact of operation types, the state-of-the-art work employs 5 operation types in Fig. 2(a), while the investigated one employs 4 operation types (i.e., without Zeroize). Details of the experimental setup are provided in Section 4. The experimental results are presented in Fig. 2(b), from which we draw the following key observations.
@@ 1.2 Case Study and Research Challenges
  • The solution should incorporate the memory constraints into its searching algorithm to ensure the applicability and efficiency of the SNN deployments on the targeted hardware.
1.3 Our Novel Contributions
Towards this, we propose SpikeNAS, a novel NAS framework that can quickly find the appropriate SNN architecture with high accuracy under the given memory budgets from the hardware platforms. It employs several key steps as the following (an overview is in Fig. 3 and the details in Fig. 4).
  • Analysis of the Impacts of Network Operations (Section 3.1): It aims to understand the significance of each pre-defined operation type, i.e., by removing the investigated operation type from the search space and then observing the output accuracy.
  • Network Architecture Enhancements (Section 3.2): It aims to maintain or improve accuracy of the network architecture, while reducing the searching time. It is done by optimizing the cell operation types and the number of cells in the network architectures.
  • A Fast Memory-aware Search Algorithm (Section 3.3): It aims to find an SNN architecture that achieves high accuracy and meets the memory budget, by performing search for each individual cell, while monitoring the memory cost of the investigated architecture.
  • SNN Quantization (Section 3.4): It further reduces the memory footprint of the generated SNN model through the post-training quantization (PTQ) technique, while considering the accepted accuracy degradation.
Key Results: We evaluate our SpikeNAS framework using a Python-based implementation on an Nvidia RTX A6000 GPU machine. The experimental results show that our SpikeNAS improves the searching time and maintains high accuracy compared to state-of-the-art while meeting the memory budgets, e.g., our SpikeNAS improves the searching time by 29x, 117x, and 3.7x for CIFAR10, CIFAR100, and TinyImageNet200, respectively.
@@ 1.3 Our Novel Contributions
Figure 4: Overview of our SpikeNAS framework, and its key steps for generating an SNN architecture that offers high accuracy while meeting the given memory constraint from an embedded AI system (e.g., autonomous mobile agent). The novel contributions are highlighted in blue boxes.
2 Preliminaries
2.1 Spiking Neural Networks (SNNs)
SNNs are considered the third generation of NN computation model [17]. SNNs typically employ a LIF neuron model due to their bio-plausible data representation and computation [22][30][42, 43, 44, 45], whose neuronal dynamics can be formulated as follows.
@@ 2.1 Spiking Neural Networks (SNNs)
2.2 Neural Architecture Search (NAS) for SNNs
NAS techniques for SNNs [5][28][39, 40, 41][46] can be classified into two approaches, i.e., “NAS with training” and “NAS without training”.
NAS with Training: In this category, the prominent work is the AutoSNN [5], which aims at finding an SNN architecture with high accuracy and low number of spikes. It trains a super-network that includes all of the possible candidates of SNN architectures, then performs a search to find a subset of network architecture that achieves the target. However, this NAS approach requires a huge searching time as well as a huge memory and energy consumption, as it needs to perform training for a large network before searching and determining the desired SNN architecture.
NAS without Training: Initially, NAS without training approach was proposed in ANN domain [47], and then leveraged for SNN domain [28]. In this category, the prominent work is the SNASNet [28], whose key idea is to quickly measure the capability of a network to reach high accuracy at early stage, thereby avoiding the training phase to evaluate the network. It is motivated by the observation that an architecture with distinctive representations across different samples is likely to achieve high accuracy after training [47]. To measure the representation capability of a network, the SNASNet employs the following key steps (our SpikeNAS framework also follows the same steps for evaluating the network).
@@ 2.2 Neural Architecture Search (NAS) for SNNs
To achieve this, SNASNet explores the architecture candidates of the cell (i.e., a directed acyclic graph that represents network topology and operations); see Fig. 2(a). Afterward, it evaluates the score of the investigated architecture, then trains the architecture with the highest score [28]. However, it has several limitations as follows.
  • It has a fixed number of cells (i.e., 2 cells) as shown in Fig. 5 and it applies the same architecture for all cells, thus reducing the possibility to find a better SNN architecture (e.g., higher accuracy).
2.3 Neural Cell-based Strategy for NAS
Neural cell (or simply cell)-based strategy has been widely used for NAS in ANNs. The motivation of cell-based NAS is to provide a unified benchmark for NAS algorithms [38]. Its approach is to search for the connection topology and the corresponding operations inside each cell; see Fig. 2(a). A cell typically consists of 4 nodes and 5 pre-defined operation types: No-Connection (Zeroize), Skip-Connection (SkipCon), 1x1 Convolution (1x1Conv), 3x3 Convolution (3x3Conv), and 3x3 Average Pooling (3x3AvgPool). The connection between 2 nodes employs a specific pre-defined operation type. Therefore, there are 15625 architecture candidates in each cell. The state-of-the-art work (i.e., SNASNet) employs a 2-cell SNN architecture for its NAS technique; see its macro-architecture in Fig. 5. In this work, we also employ the same macro-architecture for our SpikeNAS framework, and focus on the feedforward topology as it has been widely used in the SNN community.
3 SpikeNAS Framework
To address the above-discussed research problem and challenges, we propose a novel SpikeNAS framework; see the detailed flow in Fig. 4. Its key steps are discussed in the Section 3.1 - Section 3.3, consecutively.
3.1 Analysis of the Impact of Operation Types
Discussion in Section 1.2 suggests that employing all pre-defined operations might not lead to optimal solution. Therefore, to understand the significance of each pre-defined operation, we perform an experimental case study. Here, we remove a specific operation type, then perform NAS considering a macro-architecture in Fig. 5 with 5000x search iteration. Since we employ the NAS without training approach, then this step is performed before the search process and the training phase. Experimental results are shown in Fig. 6, from which we draw the following key observations.
@@ 3.1 Analysis of the Impact of Operation Types
    Removing the 3x3Conv operation from the cells significantly reduces the accuracy, as this operation employs 3x3 kernel size for extracting important features from input feature maps. Hence, removing the 3x3Conv make the network difficult to perform learning properly. Therefore, the 3x3Conv operation should be kept in the search space.
@@
3.2 Network Architecture Enhancements
3.2.1 Optimization of the Operation Types in the Neural Cells
In Section 3.1, we observe that some operations can be removed from the search space to improve the searching time without sacrificing accuracy. Hence, we optimize the number of operation types based on their significance on the accuracy. We propose two variants of operation sets that can be considered in the search space to provide multiple optimization options (see an overview in Fig. 7), as follows.
@@ 3.2.1 Optimization of the Operation Types in the Neural Cells
3.2.2 SNN Macro-Architecture Adjustments
@@
  • 1-Cell Architecture: It aims at reducing the memory footprint, but might reduce the accuracy as compared to the 2-cell architecture; see an illustration in Fig. 8. Its key idea is removing one cell from the 2-cell architecture; see an overview in Fig. 9(a).
  • 3-Cell Architecture: It aims at improving the accuracy, but might incur a larger memory footprint as compared to the 2-cell architecture; see an illustration in Fig. 8. Its key idea is adding one cell to the 2-cell architecture; see an overview in Fig. 9(b).
@@
3.3 A Fast Memory-aware Search Algorithm